the base of a 20 foot ladder is 15 ft from the house. how far up the house does the ladder reach
Answer by rothauserc(4718)

let h be how far up the house the ladder reaches
:
15^2 + h^2 = 20^2
:
225 + h^2 = 400
:
h^2 = 175
:
h = 13.23
:
*************************************************
the ladder reaches 13.23 feet up the house
*************************************************
: